# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 10 Uncertainties in the MB Calibration for SNe Ia
A current and major uncertainty in the SN Ia distance scale is the relatively small number of SNe Ia calibrators in the local universe that are accessible to HST where Cepheid, TRGB or JAGB distances can be measured. At present, there are 42 SNe Ia located in 37 galaxies where Cepheids have been discovered by the SHoES team. In Figure 15 we plot the absolute peak B magnitudes, MB, for this sample of 42 SNe Ia as a function of distance modulus, all based on the Cepheid distance scale and SN Ia magnitudes from R22. All of the data points (blue and orange) are from the R22 data set. The orange points are the subset of SNe Ia observed as part of this study. The JWST sample selection was based on proximity, to minimize any possible systematics due to crowding/blending effects.
@@
A 3 trend can be seen in this figure, in the sense that the supernova absolute magnitudes become fainter with increasing distance. In addition, for the nearest third of the SN Ia sample, the average MB magnitude is -0.07 mag brighter than the remaining more distant sample, or 2.8% in distance. The nearby sample may be fainter because it is not sampling the full distribution, given the small-number statistics. Alternatively, the lower signal-to-noise and poorer resolution of the more distant sample may be contributing a systematic uncertainty. (And both effects may also be at play.) However, it is well to keep in mind that at distances greater than 40 Mpc, the Cepheids are crowded by stars within an area 33 times larger than at 7 Mpc, while, simultaneously the signal-to-noise is lower for the more distant sample. (At distances greater than 20 Mpc, Cepheids are crowded by stars contained within an area at least 8 times larger than the equivalent resolution element (fitting radius) at 7 Mpc.) Future data will be required to assess the relative contributions of these two possibilities.
At this juncture in our CCHP program, we have a calibrating sample of 11 SNe Ia, which will continue to be augmented. For context, this sample can be compared with the early SHoES samples, where, for instance Riess et al. (2009) had a total of six nearby galaxies, for which they determined a value of = 74.2 3.6 km s-1 Mpc-1, a 4.9% measurement. Riess et al. (2011) then increased the sample by two additional SNe Ia, bringing their total to eight, and resulting in a value of = 73.8 2.4 km s-1 Mpc-1, a 3.3% uncertainty. The addition of these two supernovae increased the dispersion in the SNe Ia peak MB magnitudes from 0.08 to 0.12 mag and the mean absolute magnitude dropped significantly, from -19.34 mag down to -19.27 mag, a shift of 0.07 mag. On its own, this change would decrease the distance scale by 3.3% and increase the Hubble constant by a similar percentage. However, other improvements to the SHoES analysis led to a change in the opposite direction so that decreased by 0.8%. These results agreed extremely well with their current estimate of = 73.04 1.04 km s-1 Mpc-1, based on a sample of 42 SNe Ia.
@@
Is the existing calibrating SNe Ia sample accurate enough to provide an estimate of at the 1% level? Some of the challenges in measuring accurate distances to the nearest galaxies in this sample can be seen in the lower panel of Figure 15. Shown is a comparison of the MB magnitudes for SNe Ia calibrator galaxies in common to the R16 and R22 studies (based on distances published in their Tables 5 and 6, respectively). The HST data for these original 19 galaxies (20 SNe Ia) were first analyzed in R16, and the same raw data were then re-analyzed in R22. Differences up to 0.3 magnitudes in distance modulus were measured, with a net increase in distance modulus of 0.033 0.024 mag, leading to an increase in MB (R22-R16) of 0.037 0.033 mag in the mean. Considered in isolation, this difference alone would result in a decrease in of 1.7% (but other improvements elsewhere resulted in no net change to ). However, most of the (most challenging distant) galaxies beyond 40 Mpc have not yet been imaged at the improved resolution of JWST.
@@
The augmented R22 sample contains 42 SNe Ia, and its statistical uncertainty is smaller than available for our current JWST sample. However, it must be noted that 60% of the R22 sample galaxies are located at distances 20 Mpc and 25% are more than 40 Mpc distant. The more distant galaxies, on average, have lower signal to noise (S/N) and decreased resolution. Future high-resolution data will be needed to both improve the statistical precision of those distances and to rule out any further potential systematic effects.
In the left-hand panel of Figure 16, we compare the MB probability density functions for the SNe Ia in the R22 sample, those with combined SN Ia peak mB and Cepheid distance modulus errors ( ) less than 0.15 mag (23/42 objects). In blue are the 11 SNe Ia that are located closer than 31.8 mag in distance modulus (d 23 Mpc) ; in red are the 12 SNe Ia more distant than 31.8 mag. For this highest-quality (in both SNe Ia light curve and Cepheid distance precision) sample, there is again a clear systematic offset with distance in the peak SN Ia brightness, in the sense that the nearby sample is brighter in the mean. The right-hand panel shows the PDFs for the sample of SNe Ia with larger uncertainties ( 0.15). In this case, the peak brightness is fainter for both the nearby and more distant SNe Ia. These comparisons highlight once again the critical role that this intermediate rung is playing in the SNe Ia distance scale, and the need for additional JWST data and a larger SNe Ia calibrating sample.
@@ 11 Error Budget
The statistical and systematic uncertainties for each of the three distance indicators are discussed in detail in the companion papers of O24, H24 and L24. Here we note that in the determination of , the overall uncertainties for the individual galaxy distances become statistical uncertainties. For example, while the aperture corrections or reddening corrections for an individual galaxy contribute to its systematic uncertainty, for an ensemble of galaxies, the aperture corrections or reddening corrections become a source of random/statistical uncertainty. The overall statistical uncertainty is determined by the number of host galaxy calibrators, and the total number of SNe Ia contained therein.
In Table 5 we summarize the sources of the statistical and systematic uncertainties in our measurement. We list the statistical and systematic errors separately. The combined statistical errors are given by: . (The individual statistical uncertainties were discussed in §9.4.) The individual systematic errors, as discussed in detail in O24, H24, and L24 are 3.10, 1.54 and 1.90 km s-1 Mpc-1, for Cepheids, TRGB and the JAGB, respectively. The systematic errors common to all three methods are that for the zero point (anchored to NGC 4258), and the photometric zero point; otherwise the uncertainties are independent. A description of the components of the systematic errors is given in the notes to Table 5. Based on our Bayesian analysis with a flat prior, as described in §9.4 and illustrated in Figure 11, our adopted value of the Hubble constant and its uncertainty, as derived from our new JWST distances applied to the CSP sample of SNe Ia is = 69.96 1.05 (stat) 1.12 (sys) km s-1 Mpc-1.
Source of Error | Random Error | Systematic Error | Reference |
TRGB zero point | 2.5% | 2.2% aafootnotemark: | H24, §9.4 |
JAGB zero point | 2.7% | 2.8% bbfootnotemark: | L24, §9.4 |
Cepheid zero point | 2.6% | 4.3% ccfootnotemark: | O24, §9.4 |
Combined methods ddfootnotemark: | 1.5% | 1.6% | §9.4 |
@@ 11.1 Overall Systematic Uncertainties
Combining the results for the three different methods provides a means of averaging over the systematic effects that independently affect any given method. There are some systematics, however, that are common to all of the methods.
@@
As described in §10 and illustrated in Figures 15 and 16, a current uncertainty in the SN Ia distance scale is the relatively small number of calibrators with well-observed SNe Ia, and distances measured with Cepheid TRGB or JAGB distances. A concern is the trend that SN Ia absolute magnitudes get fainter with increasing distance (and increasing errors), a result that could be due to statistical effects and/or systematic effects in the measured distances, particularly in galaxies where crowding/blending effects could be more severe.
@@
The calibrations of the TRGB, JAGB stars and Cepheid Leavitt law for the JWST data presented in this paper are all based on the geometric distance to one galaxy, NGC 4258, and the three methods therefore share any systematic errors (known or unknown) in that determination. The currently cited total uncertainty in the NGC 4258 distance is 1.5% (Reid et al., 2019). Unfortunately, at 7.5 Mpc (and with a central accretion disk that is highly inclined, allowing accurate measurements of its maser orbits), NGC 4258 is the only nearby galaxy that is suitable for the calibration of the TRGB, JAGB stars and Cepheids based on masers. (The next nearest galaxy is at a distance of over 50 Mpc, with a measured uncertainty of 9% (Pesce et al., 2020).) With NGC 4258 alone, it is not possible to determine any underlying galaxy-to-galaxy differences inherent in the maser method, or to assess the impact of any systematics tied to assumptions in the modeling. In forthcoming papers, we will continue to improve upon and apply additional zero-point anchors to each of the three distance determination methods; for this study, we have adopted a consistent set of distances and calibration based on JWST data alone. We have explicitly tested the NGC 4258 calibration in the case of the TRGB (§12.3), finding excellent agreement with previous studies that were based on a calibration with four geometric anchors.
@@
The measurement of the Cepheid distance to NGC 4258 carries with it the uncertainty in , the slope of the sensitivity of the Leavitt law to metallicity, the value of which is still a subject of active debate in the literature. For the purposes of this study, we have adopted a value of = -0.2 mag/dex (as adopted by R22); this translates to less than a 1% (i.e., 0.7%) metallicity correction to , in the sense that decreases when the correction is applied.
However, if = -0.5 mag/dex as some Gaia studies have concluded (e.g., Ripepi et al., 2020), it would be a 2% correction. This becomes a more serious issue when using the LMC as an anchor galaxy. R22 adopt [O/H] = -0.29 for the LMC; and this would translate to a 6.3% metallicity correction for . Furthermore, the LMC’s geometric distance currently has the highest precision of all of the current anchors, making it the largest contributor to the estimation of in the multi-anchor analyses. Future work is needed to establish, with confidence, the metallicity sensitivity of the Leavitt law. In combining the three methods, an error in the Cepheid metallicity correction will enter as part of the overall systematic uncertainty in our value of .
The impact of metallicity on the Cepheid calibration presented here is small because the average metallicity of our sample is close to that of NGC 4258 itself, based on the metallicity scale inferred from HII regions adopted by R22. However, it is worth emphasizing that there remains a trade-off between having only a single anchor galaxy with a small metallicity correction versus the case of more anchor galaxies, each being impacted by increasingly uncertain metallicity corrections.
@@
For all three applications, we have adopted the same reddening law, which is assumed to be universal, with the same ratio of total-to-selective absorption, . In addition, for the JAGB and TRGB, where foreground reddening corrections are required (for the Cepheids a total reddening correction is measured), there is an uncertainty due to the zero-point calibration of the extinction law. Schlegel et al. (1998) estimates this uncertainty to be = 0.02 mag, which corresponds to an extinction uncertainty of AJ = 0.016 mag (0.7%). In an era of accurate cosmology, this level of uncertainty becomes more significant.
@@
In this analysis, we use SNe Ia to extend the distance range for the determination. Any remaining systematic error intrinsic to SNe Ia will be shared by all three of the methods presented here. These systematics might include inconsistent calibrations of photometric zero points across different surveys and instruments that can lead to systematic offsets (e.g., in the case of Pantheon+); in addition the calibrator light curves are generally not part of the well-calibrated CSP survey, which makes up the Hubble flow sample. Further uncertainties may arise in the corrections for dust extinction (e.g., existence of more than one type of dust: differences in the dust surrounding the SN Ia, variations in the extinction law, etc.), intrinsic scatter in the peak luminosity of SNe Ia, and/or effects such as the host galaxy mass step, which may not be fully constrained, potentially leading to residual, uncorrected biases in linking the calibrator SN Ia luminosities to the more distant SNe Ia.
6.
Finally, the results of an MCMC analysis may not be as robust as the confidence implied by the likelihood analysis. These include the fact that the probability might be accurately treated as a Gaussian near the peak, but it may not decrease as fast as a Gaussian in the tails; there may be unknown systematic errors revealed as the statistical uncertainty decreases and a systematic floor is reached; and the likelihood function may have uncertainties and/or assumptions that have not been included in the quoted uncertainty (e.g., Liddle, 2009; Hogg & Foreman-Mackey, 2018).
We note that the uncertainties that we are quoting are larger than those given, for example, by R22. In summary there are several reasons for this difference. (1) Our sample of SN Ia calibrators is smaller by a factor of four than that of R22, resulting in a factor of two larger statistical uncertainty. This uncertainty, however, is at least partially offset by the fact that we have three distance methods. (2) We have only a single anchor galaxy, NGC 4258, whereas R22 have three anchors (LMC, the Milky Way and NGC 4258). However, as discussed in §12.3, there is excellent agreement with the -band calibration of the TRGB with four anchors. (3) We have adopted a more conservative estimate of the uncertainty in the metallicity dependence of the Leavitt law, reflecting the current uncertainty in the Gaia absolute calibration, as well as the range in published values of the metallicity slope parameter, (see §11.1). (4) We separate our statistical and systematic uncertainties and do not marginalize over the uncertainties.

### v2

@@ 11 Uncertainties in the MB Calibration for SNe Ia
Achieving a 1% accuracy in measurements presents considerable challenges. Several concerns within the current MB calibration must be addressed before the true uncertainty in the local value can be confirmed. A significant uncertainty stems from the limited number of local supernovae accessible to HST and JWST for Cepheid, TRGB, or JAGB distance measurements. Here we examine several issues impacting the accuracy of the current value derived from the calibration of SNe Ia, and assess the suitability of the existing calibrator sample for a 1% accuracy estimate. Our conclusion is that claims of such precision are currently premature.
1) At present, there are 42 SNe Ia located in 37 galaxies where Cepheids have been discovered with HST by the SHoES team. For the closest SNe Ia (23 Mpc) sample with the smallest errors ( 0.15 mag) the weighted mean MB magnitude is 0.08 mag (3.8%) brighter than that for the rest of the sample (-19.32 0.03 versus -19.24 0.03 mag). While the samples are small and the difference amounts to approximately 2, this difference is large enough to warrant future study to determine if it is due to random statistical variation or an underlying systematic effect. Accurately constraining the systematic uncertainties is crucial. Even a subtle overall 0.03 mag difference corresponds to 1 km s-1 Mpc-1, the entire size of the SHoES published uncertainty. Therefore, increasing the sample size and improving the statistics remain a key area for future follow-up.
As an example of how potential systematics can impact the measurement of Cepheid distances, in a detailed recent analysis of the period distributions of Cepheids in the SHoES sample, Högås & Mörtsell (2024) found that anchor galaxies exhibit a shorter period range compared to SN Ia host galaxies. Furthermore, galaxies with shorter period ranges, more closely resembling the anchors, yielded significantly lower values. They note that this effect is consistent with a correlation between SN Ia absolute magnitude and distance modulus shown in Appendix B since shorter-period (fainter) Cepheids aresystematically undiscovered because of photometric incompletenessin galaxies at increasingly larger distances.
@@
2) The lower signal-to-noise, as well as the poorer resolution of the more distant sample may be contributing a systematic uncertainty. For a desired total uncertainty of 1%, it is well to keep in mind that at distances greater than 40 Mpc, the Cepheids are crowded by stars within an area 33 times larger than at 7 Mpc (the distance to the calibrating galaxy, NGC 4258) while, simultaneously the signal-to-noise is lower for the more distant sample. (At distances greater than 20 Mpc, Cepheids are crowded by stars contained within an area at least 8 times larger than the equivalent resolution element (fitting radius) at 7 Mpc.) Future data will be required to assess whether the more distant sample of Cepheid galaxies are free from systematic crowding bias or whether lower signal-to-noise results in increased uncertainties in the photometry of more distant galaxies.
R24 have argued that crowding effects have been ruled out at an 8.2- level of confidence. However, of the five galaxies for which they compare HST and JWST distance moduli, one is extremely nearby (NGC 5643 at 13 Mpc), two (NGC 1448 and NGC 5584) showed the smallest differences between their 2016 and 2022 distance moduli (0.01 - 0.02 magnitudes) and thus are the least likely to have significant crowding issues, and only one out of the five galaxies, NGC 5468, is at a distance of 40 Mpc. R24 assert that at 40 Mpc one would expect a 0.30 mag effect to explain the tension due to crowding. (This of course would hold only if all of the tension were to be explained by crowding, a highly unlikely explanation.) However, it should be noted that a 0.03 mag systematic error (i.e., comparable to what was found in the nearby sample) would result in a systematic error in of 1 km s-1 Mpc-1, the entire size of the R22 uncertainty. As noted, most of the (most challenging distant) galaxies beyond 40 Mpc have not yet been imaged at the improved resolution of JWST.
@@
3) Some of the challenges in measuring accurate distances to the nearest galaxies in this sample can be seen in the Figure 15. Shown is a comparison of the distance moduli for SNe Ia calibrator galaxies in common to the R16 and R22 studies (based on distances published in their Tables 5 and 6, respectively). The HST data for these original 19 galaxies (20 SNe Ia) were first analyzed in R16, and the same raw data were then re-analyzed and re-calibrated in R22. Differences up to 0.3 magnitudes in distance modulus were measured, with a net increase in the mean distance modulus (R22-R16) of 0.035 0.025 mag (and a brightening of MB by the same amount). Considered in isolation, this systematic difference would result in a decrease in by 1.6%. Indeed, if one retroactively corrects the R16 distances to the R22 scale correcting for the 0.035 mag offset, the SHoES value becomes 72.05 1.74 km s-1 Mpc-1, resulting in a lower (2.6) and much less interesting Hubble tension (see also §12).
@@
It must be noted that 60% of the R22 sample galaxies are located at distances 20 Mpc and 25% are more than 40 Mpc distant. The more distant galaxies, on average, have lower signal to noise (S/N) and decreased resolution. Fortunately, this is an issue that can be resolved with further higher resolution JWST data for the more distant sample, and it is clearly warranted before crowding can be definitively ruled out as a potential systematic.
4) The photometric calibration of SNe Ia is still undergoing changes with systematic shifts the size of the SHoES total error bar. In Figure 16, we show the change in adopted apparent SN Ia magnitudes (mb) between the Riess et al. (2016) and Riess et al. (2022) studies plotted as a function of MB. The point sizes scale with the weight, w, for each object in the MB calibration, as defined by
The shift in apparent magnitude is systematic: two thirds (13 of the 19) of the SNe Ia have fainter magnitudes in the R22 study. The mean shift in apparent magnitude amounts to 0.03 mag, partially compensating for the decrease in that would result from the increase in the SHoES mean distance moduli between R16 and R22 (see §9.1).
5) The faintest supernova in the Riess et al. (2022) sample is SN2007af in the galaxy NGC 5584. From Figure 13, this supernova can be seen at the faint end of the distribution. For the nearest half of the sample (comparable in distance to our JWST sample), SN 2007af is the most discrepant point in the weighted average. In this nearest half of the sample, the weighted mean = -19.284 0.039 mag (error of the mean). Removing the single supernova, SN 2007af increases the weighted mean magnitude to MB = -19.316 0.032 mag. (This value can be compared to the value of MB = -19.317 0.033 mag that is obtained for the nearby JWST sample using the SHoES data.)
In comparing 17 SNe Ia in common between the SHoES and CCHP studies, H25 notes how the pull of outliers is significantly reduced in the TRGB calibration. For this overlapping sample, recall that in the case of the SHoES analysis, if SN2007af is dropped from the sample, MB shifts by 0.029 mag (1.3%), corresponding to a shift in of -1.02 km s-1 Mpc-1. Note that this shift, due to the exclusion of a single SN Ia calibrator, is as large as the entire R22 uncertainty in . However, this is not the case for the CCHP TRGB calibration, where MB shifts by a negligible amount of only 0.004 mag, or -0.15 km s-1 Mpc-1.
In the case of the total sample of 42 SNe Ia, dropping SN2007af results in a decrease in of 0.5 km s-1 Mpc-1 (half of the R22 uncertainty) and it lowers the dispersion in the weighted mean from 0.129 to 0.115 mag. We emphasize again that a major advantage of the TRGB distance scale is the ability to measure stars in the outer regions of galaxies where dust and crowding effects are minimal. The errors in the TRGB distance moduli range from 0.04 to 0.10 mag (a factor of 2.5), whereas for Cepheids, the uncertainties differ by an order of magnitude, ranging from 0.037 to 0.253 mag (a factor of 6.8); i.e., spanning nearly a factor of 50 in the variance (and therefore strongly weighting different supernova disproportionately, with some being 50 times more influential than others) in the calibration.
6) H25 discusses at length the CCHP TRGB and SHoES Cepheid calibrations of the CSP-I + CSP-II, as well as SuperCal (Scolnic et al., 2015) and Pantheon+ (Scolnic et al., 2023) SN Ia data. Restricting the comparison of CSP and Pantheon+ to the same overlapping subset of SNe Ia (22 in the case of CSP-I and 14 for SuperCal, H25 finds that in the case of the CSP data, the larger HST + JWST sample retains a stable central value when compared to the JWST sample only, differing by only -0.11 km s-1 Mpc-1. However, this does not occur in the case of the SuperCal and Pantheon+ data. In the latter case, the difference in between the SuperCal and Pantheon+ data is +1.20 km s-1 Mpc-1. Note that this difference cannot be explained by differences in the TRGB and Cepheid distances (which now agree to , see §9.1). This difference, once again greater than the total R22 quoted uncertainty, is now largely due to changes in the apparent magnitudes, mB, introduced in the transition from SuperCal to Pantheon+ (e.g., see Figure 16). We note also that six of the eight calibrations discussed by H25 yield values of 70 km s-1 Mpc-1. A single calibration gives 71 km s-1 Mpc-1. The Pantheon+ calibration is a significant outlier in the HST + JWST TRGB calibration, shifting the Hubble constant by +1.7 km s-1 Mpc-1.
7) Half of the weight in the SHoES sample of MB values comes from only a dozen SNe Ia. These 12 SNe Ia represent only 29% of the total sample of 42. An already small sample of 42 SNe Ia is not providing statistical (increased sample size) gains. An effective sample size can be computed as follows:
That is, the effective size of the R22 Cepheid sample is equivalent to only 31 SNe Ia, or 74% of the total sample. (The effective size of the CCHP TRGB sample is 21 compared to a total of 24, or 88%; i.e., the weights are more evenly distributed.) These limitations, however, underscore the need for a larger sample of nearby SNe Ia.
In summary, recent changes to the Cepheid distance scale are illustrative of the types of potential (at-the-time-unknown) systematic errors, having sizes comparable to the total current estimated uncertainties. Furthermore, recent revisions to the originally published apparent magnitudes of these supernovae, comparable in scale to the entire quoted SHoES uncertainty, suggest that the SN Ia calibration may not have yet converged. Despite the SHoES program’s use of 42 SNe Ia calibrated with Cepheid distances, the current Cepheid distance and SN Ia calibration uncertainties result in less than one-third of the supernova sample (12 objects) contributing half the overall calibration weight. The implications for a claim for physics beyond the standard model clearly demand a larger sample of nearby SNe Ia on which to base the calibration. Even a subtle difference, such as a 0.03 magnitude variation, can significantly impact the value of at a level of the current quoted error budget. Below we review the tension in , and conclude that the current state of the nearby calibration does not support a tension level of 5.
12 The Tension
Given that significant systematic offsets are still being identified empirically that are comparable to the current total estimated 1 uncertainty indicates that representative or realistic uncertainties have not been previously accounted for in estimating the true tension. The claim of 5 tension does not allow for the observed systematic evolution of the measurements, which violates the assumption of statistical independence and likely inflates the significance.
A tension of 5 follows from absorbing the uncertainties in into a single quantity as follows:
where 1.04 and 0.5 km s-1 Mpc-1.
As an illustrative example, given here to evaluate the impact of these shifts on the calculated Hubble tension, we consider several recent changes to the SHoES Cepheid distances. We compare the distances in the (baseline) R22 Cepheid sample with those obtained three years previously in the Riess et al. (2019) [hereaftr, R19] study. R19 did not tabulate their distance moduli, but the average difference between the R22 and R19 Cepheid distance moduli amounts to mag (which is almost a 3% change expressed in distance). This difference reflects:
1.
The difference between R19 and R16, where R16 values were updated as follows:
An update to the LMC distance from Pietrzyński et al. (2013) to Pietrzyński (2019) and the recalibration of HST/WFC3-IR LMC Cepheid photometry amounting to a total of R19-R16 = -0.029 mag. This update resulted in a value of = 74.22 1.82 km s-1 Mpc-1 based on the LMC alone (differing from that of = 72.04 2.67 in R16). Combining the LMC, Milky Way and NGC 4258 distances resulted in a value of = 74.03 1.42 (corresponding to a difference of only +0.006 mag from the LMC alone).
- 2.
The average difference between R22 and R19 thus amounts to:
[(R22-R16) - (R19-R16)] = [(+0.035) - (-0.029 + 0.006)] = +0.058 mag.
We first subtract the systematic shift in (2.00 km s-1 Mpc-1) due to the revision to the R19 Cepheid distances (i.e., the impact on due to the difference of +0.058 mag in the average distance moduli). We take the recent concordance in the SHoES Cepheid and CCHP TRGB distances as an indication that this correction is reasonably robust. Given the fact that the SN Ia calibration has not yet converged among the various projects (see also DESI Collaboration et al., 2024), we then take the difference in the apparent SN Ia magnitudes (the shift of 0.03 mag) as an indication of outstanding systematic uncertainties in the SN Ia calibration (and also reflective of the difference in the Pantheon+ and CSP plus SuperCal (see H25) SN Ia data sets). For the purpose of this exercise, we retain the small 1% uncertainty quoted by R22 and add in quadrature the uncertainty in the SN Ia calibration. We then recalculate the tension as follows:
where = 74.03 - 2.00 = 72.03 km s-1 Mpc-1; = 1.00 in km s-1 Mpc-1, and = = 1.44.
When including this systematic uncertainty in the supernova calibration, the tension is reduced to 3.0 , and is reflecting a more conservative assessment of the measurement’s reliability. Note also that this simple calculation assumes an uncertainty of = 1.04, which may still be an underestimate. (Adopting the error on the R19 value of = 1.42 results in a tension of only 2.6.) In their discussion of additional potential, unaccounted-for uncertainties, Högås & Mörtsell (2024) independently conclude that differences in the period distributions of the anchor relative to host galaxies results in only a 2.4 tension (when using the Milky Way zero point calibration).
We conclude that considerably more caution is warranted before concluding that the Hubble tension is at the level of 5. A 5 tension implies that there is only a 0.00003% probability (a vanishingly small number where astrophysical measurements are concerned) that the observed difference is due to random fluctuations. We note also that a 5 tension is only meaningful for a normal distribution. Claiming that the CDM model is incomplete carries with it the implicit assumption that it is the local value of that is the correct value, holding the standard model as the one in need of change. However, the tension itself carries no information on which value is in need of updating. We have already seen how the SHoES distance scale has shifted by nearly 1.6%, and the SN Ia magnitudes have shifted by 1.4%, both systematic differences comparable in size to the R22 total 1.4% quoted uncertainty. While it is not possible to assess the size of unknown systematics in advance, it is possible to examine recent prior estimates with an eye to previously hidden systematic uncertainties. While differences of order the size of the total quoted uncertainty are still at play, a claim of 5 tension may be premature.
@@ 13 Error Budget
13.1 Distance Errors
The statistical and systematic uncertainties for each of the three distance indicators are discussed in detail in the companion papers of H25 and L25. Here we note that in the determination of , the overall uncertainties for the individual galaxy distances become statistical uncertainties. For example, while the aperture corrections and reddening corrections for an individual galaxy contribute to its systematic uncertainty, for an ensemble of galaxies, the aperture corrections and/or reddening corrections are uncorrelated from galaxy to galaxy, and become a source of random/statistical uncertainty. The overall statistical uncertainty is determined by the number of host galaxy calibrators, and the total number of SNe Ia contained therein.
In Table 6 we summarize the sources of the statistical and systematic uncertainties in our measurement. We list the statistical and systematic errors separately. The individual statistical uncertainties were discussed in §10.4.) The systematic error common to both the TRGB and JAGB methods is that for the zero point (anchored to NGC 4258). As discussed in §8, there is an independent systematic uncertainty for each method that results from the measurement uncertainty in the TRGB or JAGB magnitude for NGC 4258. These two errors are added in quadrature and given in Table 6. Finally, we introduce an additional 1% error term, denoted as , to account for residual uncertainties in the supernova calibration, including photometric calibration of MB and peculiar motion effects.
13.2 Supernova Errors
For SN Ia observations, a common limitation of all measurements (SHoES/Pantheon+ and CCHP/ CSP alike) is that they rely on data collected using various telescopes, instruments, and filter combinations, making the sample of nearby host SN Ia galaxies heterogeneous. However, for distant SNe Ia, the CSP has an advantage, as all observations were conducted at Las Campanas using consistent filter combinations and detectors. In some cases CSP data were obtained for both nearby and distant SN Ia samples, in which case the error in zero point cancels. However, since the CSP did not begin until 2004, any SNe Ia discovered earlier lack CSP observations. A similar limitation applies to SNe Ia in northern hemisphere galaxies that were inaccessible to southern telescopes at Las Campanas (e.g., 2011fe in M101). The typical error in optical zero points is around 0.01 mag; however, to be conservative, we adopt a systematic uncertainty of 0.015 mag to account for the inhomogeneous nature of the current calibration.
All of the nuisance parameters in the SN Ia magnitudes, as characterized in Equation 1, are parameters in the MCMC analysis, and are included in the final uncertainty in from that analysis. The CSP K-corrections are all computed with the same time series spectral energy distribution template, which could result in a systematic error. However, as can be seen from Lu et al. (2023), Figure 16, the systematic uncertainties in the near-infrared K-corrections are, in any case, very small (1%); the optical data have more spectra contributing to the template, and their uncertainties are also very small.
13.3 Adopted Uncertainty on H0
We adopt the uncertainties as described above in §13 and shown in Table 6. Based on our HST + JWST TRGB sample of 24 SNe Ia, as described in §10.4 and illustrated in Figure 10, our adopted value of the Hubble constant and its uncertainty, as derived from our new JWST distances applied to the CSP sample of SNe Ia is = 70.39 1.22 (stat) 1.33 (sys) 0.70 [SN] km s-1 Mpc-1.
| Source of Error | Random Error | Systematic Error | Reference | |
| TRGB (JWST) | 2.6% | 1.9%aafootnotemark: | 1% | H25, §14.1,§10.4, this section |
| JAGB (JWST) | 3.0% | 2.4%bbfootnotemark: | 1% | L25, §14.1, §10.4, this section |
| Combined distances | 2.4% ccfootnotemark: | 1.8% | 1% | §14.1,§10.4 |
| TRGB (HST + JWST) | 1.7% | 1.9% | 1% | §10.4, this section |
@@ 13.4 Overall Systematic Uncertainties
We describe here the systematic effects that are common to the local distance (TRGB, JAGB, Cepheid and SN Ia) scales. Although generally small, in an era of accurate cosmology, percent-level uncertainties will become more significant.
@@
The calibrations of the JAGB stars and TRGB for the JWST data presented in this paper are both anchored to the geometric distance to one galaxy, NGC 4258, and these two methods therefore share any systematic errors (known or unknown) in that determination. The currently cited total uncertainty in the NGC 4258 distance is 1.5% (Reid et al., 2019). However, the TRGB distances measured with HST (Freedman, 2021), and tied to four geometric anchors, agree to within 1% of the JWST TRGB calibration based on NGC 4258 alone. In forthcoming papers, we will continue to improve upon and apply additional zero-point anchors to the TRGB, JAGB and Cepheid distance scales.
@@
As discussed in §11, a current key source of uncertainty in the SN Ia distance scale is the relatively small number of well-observed SNe Ia with calibrators that have distances measured using JWST. One concern is a trend (see Figure B1) in which SN Ia absolute magnitudes appear to become fainter with increasing HST-measured Cepheid distances (and increasing errors). This trend could arise from statistical effects or systematic issues in the measured distances, particularly in galaxies where crowding/blending effects, lower signal-to-noise, or differences in the period ranges covered may introduce biases.
@@
We have adopted the same reddening law, which is assumed to be universal, with the same ratio of total-to-selective absorption, . For the JAGB and TRGB, where foreground reddening corrections are required, there is an uncertainty due to the zero-point calibration of the extinction law. Schlegel et al. (1998) estimates this uncertainty to be = 0.02 mag, which corresponds to an extinction uncertainty of AJ = 0.016 mag (0.7%).
@@
In this analysis, we use SNe Ia to extend the distance range for the determination. Any remaining systematic error intrinsic to SNe Ia will be shared by all of the local distance methods. These systematics might include inconsistent calibrations of photometric zero points across different surveys and instruments that can lead to systematic offsets (e.g., in the case of Pantheon+); in addition the calibrator light curves are generally not part of the well-calibrated CSP survey, which makes up the Hubble flow sample. Further uncertainties may arise from dust extinction corrections, including the possibility of multiple dust types (e.g., in the Milky Way, the host galaxies or dust surrounding the SN Ia) and potential differences in the extinction law). Other sources of uncertainty include intrinsic scatter in SNe Ia peak luminosities and effects such as the host galaxy mass step, which may not be fully constrained. These factors could introduce residual, uncorrected biases in anchoring the luminosities of calibrator SNe Ia to those of more distant SNe Ia.
@@
Finally, the results of an MCMC analysis may not be as robust as the confidence levels implied by the likelihood analysis. These include the fact that the probability might be accurately treated as a Gaussian near the peak, but it may not decrease as fast as a Gaussian in the tails; there may be unknown systematic errors revealed as the statistical uncertainty decreases and a systematic floor is reached; and the likelihood function may have uncertainties and/or assumptions that have not been included in the quoted uncertainty (e.g., Liddle, 2009; Hogg & Foreman-Mackey, 2018).